# Observability rank conditions for analysing practical identifiability a priori

### v1

@@ Abstract
The concept of identifiability describes the possibility of inferring the parameters of a dynamic model by observing its output. It is common and useful to distinguish between structural and practical identifiability. The former property is fully determined by the model equations, while the latter is also influenced by the characteristics of the available experimental data. Structural identifiability can be determined by means of symbolic computations, which may be performed before collecting experimental data, and are hence sometimes called a priori analyses. Practical identifiability is typically assessed numerically, with methods that require simulations – and often also optimization – and are applied a posteriori. An approach to study structural local identifiability is to consider it as a particular case of observability, which is the possibility of inferring the internal state of a system from its output. Thus, both properties can be analysed jointly, by building a generalized observability matrix and computing its rank. The aim of this paper is to investigate to which extent such observability-based methods can also inform about practical identifiability. To this end, we explore a number of possible extensions of the rank tests, and discuss the purposes for which they can be informative as well as others for which they cannot.
@@ 1 Introduction
The concept of structural identifiability was introduced by Bellman and Åström (1970), who used the word “structure” to refer to the equations of a dynamic model. This property determines whether it is possible “to get insight into the internal structure of a system from input-output measurements”. When analysing structural identifiability, one imposes no restrictions on the quality and quantity of data that can be obtained from the experiments. Thus, if a certain state variable, or function, is defined as a model output, the implicit assumption is that it can be sampled at an arbitrarily high rate and with absolute precision. Since this is an ideal scenario, another concept – practical identifiability – is needed to account for the additional uncertainties caused by the limitations of real-world data, which are noisy and scarce to a varying degree (Lam et al., 2022). Structural identifiability analyses provide a binary answer: a model is either structurally identifiable or unidentifiable. In contrast, practical identifiability analyses yield a numerical answer, typically in the form of a confidence interval in a parameter value. Overviews of practical and structural identifiability have been written by Miao et al. (2011) and more recently by Wieland et al. (2021).
@@
The observability of a system describes the possibility of inferring its internal state by observing its output. The concept was introduced by R. E. Kalman in the context of linear control systems. For nonlinear systems, Hermann and Krener (1977) presented an observability test called observability rank condition; Tunali and Tarn (1987) applied a similar condition to the study of identifiability. It should be noted that, if a parameter in a model is considered as a constant state variable, (structural local) identifiability can be studied in the same way as observability. Both concepts were reviewed jointly in (Villaverde, 2019).
@@ 2.1 Structural identifiability and observability
We say that a parameter $\theta_{i}\in\theta$ of $M$ (1) is structurally locally identifiable (SLI) if there is a neighbourhood $\mathcal{N}(\theta^{*})$ such that, for any $\hat{\theta}\in\mathcal{N}(\theta^{*})$ and almost any $\theta^{*}\in\mathbb{R}^{n_{\theta}}$, $y(t,\theta^{*})=y(t,\hat{\theta})\iff\theta^{*}_{i}=\hat{\theta}_{i}$. If this relationship does not hold in any $\mathcal{N}(\theta^{*})$, $\theta_{i}$ is structurally unidentifiable (SU).
The above definition entails that, if a parameter is SLI, its value can be determined from knowledge of the output $y(t)$ and input $u(t)$ vectors over time. Structural local identifiability is similar to another property, observability. Conceptually, a state variable $x_{i}(\tau)$ is observable if it can be determined from the output $y(t)$ and any known inputs $u(t)$ of the model in the interval $t_{0}\leq\tau\leq t\leq t_{f}$, for a finite $t_{f}$. Otherwise, it is unobservable. A model is called SLI if all its parameters are SLI, otherwise it is said to be SU. Likewise, a model is called observable if all its states are observable, and unobservable if at least one of them is unobservable.
It should be noted that the properties defined above are structural but local, i.e. they tell us whether a parameter or a state variable can be uniquely estimated in a neighbourhood of (almost) any particular value, but they do not guarantee uniqueness in the whole of $\mathbb{R}^{n_{x}}$ or $\mathbb{R}^{n_{\theta}}.$ In contrast, a parameter is said to be structurally globally identifiable (SGI) if, for any $\hat{\theta}\in\mathcal{N}(\theta^{*})$ and almost any vector $\theta^{*}\in\mathbb{R}^{n_{\theta}}$, $y(t,\theta^{*})=y(t,\hat{\theta})\iff\theta^{*}_{i}=\hat{\theta}_{i}$, i.e. the neighbourhood $\mathcal{N}(\theta^{*})$ is the whole parameter space. A SGI parameter is SLI; a SLI parameter may be SGI.
@@ 2.2 A rank test for assessing observability
$M:\left\{\begin{aligned} \dot{x}&=&f\left(x,u\right),\\ y&=&g\left(x,u\right),\end{aligned}\right.$ (2)
Such a model is similar to the one given by (1), with the difference that it does not contain parameters (we will include them in the next subsection). The core idea is to build an observability matrix, $\mathcal{O}(x,u)$, that represents a map between the model output $y$ and its derivatives $\dot{y},\ddot{y},...$, on the one hand, and its state $x$ on the other. That is:
@@
where $\mathcal{L}_{f}g(x,u)$ is the Lie derivative,
$\mathcal{L}_{f}g(x,u)=\frac{\partial g(x,u)}{\partial x}f(x,u)+\sum_{j=0}^{j=\infty}\frac{\partial g(x,u)}{\partial u^{(j)}}u^{(j+1)},$
which quantifies the rate of change of a vector field along the flow described by another vector field. In this case, it describes the change of the model output function $g$ along the dynamics of the system $f$, which is equal to the derivative of the output, $\dot{y}.$ Subsequent derivatives of the output, $\ddot{y},...,y^{(n_{x}-1)},$ are calculated as higher order Lie derivatives, which can be computed recursively from the lower order ones as follows (Karlsson et al., 2012):
$\mathcal{L}_{f}^{i}g(x,u)=\frac{\partial L_{f}^{i-1}g(x,u)}{\partial x}f(x,u)+\sum_{j=0}^{j=\infty}\frac{\partial L_{f}^{i-1}g(x,u)}{\partial u^{(j)}}u^{(j+1)}$
The expressions above contain infinite summations. However, since $\dot{u}$ does not appear in the model equations (1), the summations can be truncated so that (Villaverde et al., 2019):
$\mathcal{L}_{f}g(x,u)$ $=\frac{\partial g(x,u)}{\partial x}f(x,u),$ (4)
@@ 2.4 Practical Identifiability
3 Modifying the rank test to take into account practical limitations
In this section we investigate in which ways we can use structural analyses – and more specifically the observability-based approach described in Sections 2.12.3 – to obtain information about identifiability in the face of experimental limitations. Thus, our investigation will be centred around one object, the observability-identifiability matrix $\mathcal{O}_{I}$ defined in (7), and we will discuss a number of directions in which to extend it or modify it so as to take into account ‘practical’ aspects of the estimation process.
@@ 3.1 Multiple experiments
Ovchinnikov et al. (2022) presented an algorithm that computes (i) the smallest number of experiments to maximise local identifiability, and (ii) the number of experiments to maximise global identifiability so that this number exceeds the minimal such number by at most one. This algorithm adopts a differential algebra approach.
@@ 3.2 Input characteristics
A question related to Sections 3.1 and 3.2 concerns the implementation of inputs in real applications, and how to modify the analyses to approximate them. For example, instead of applying an ideal ramp, due to experimental constraints we may be forced to apply instead a piecewise constant input. How should we treat this scenario in our analyses? One the one hand, an input consisting of a series of constant values resembles a set of multiple experiments with constant inputs. On the other hand, as the number of steps increases the input becomes closer to an actual ramp. The following example illustrates this case.
@@
Let us now perform a numerical experiment: we simulate the model to generate artificial data, applying as input a piecewise constant signal. Since our focus is on the role played by the input, we do not add any artificial noise to the output, to avoid confounding effects. The left panel of Figure 1 shows the data, the model states, and the input. Next, we try to estimate parameter $k_{12}$ by fitting the model to the data. We use a profile likelihood approach: we fix the value of $k_{12}$ and let the optimiser find the values of the remaining parameters; we repeat this procedure for different values of $k_{12}.$ The right hand panel of Figure 1 shows the resulting profile, which has a shape indicating practical identifiability and a minimum at the parameter’s nominal value ($k_{12}=7$). Hence, performing the observability rank test by approximating a piecewise constant input as a ramp yields correct SIA results (at least in this case study), while the approximation as a multi-experiment with constant inputs yields a wrong result. Intuitively, this can be understood as a failure of the multi-experiment setup to capture the continuity between experiments; a set of independent experiments does not provide all the information.
Figure 1: Model $\mathcal{M}_{3.2}$. The panel on the left shows simulations of the model input, $u(t),$ and the state variables, $x_{1}(t)$ and $x_{2}(t).$ The model output is $y(t)=x_{1}(t);$ the blue circles represent the artificial noiseless data used for parameter estimation. The panel on the right shows the resulting profile likelihood of parameter $k_{12}$. It can be seen that the parameter is structurally and practically identifiable.
@@ 3.3 Initial conditions
The definitions of structural local identifiability and observability are valid ‘almost everywhere’, that is, with the possible exception of a set of measure zero. Thus, a model that is SLI may become unidentifiable for specific initial conditions. The following model August and Papachristodoulou (2009) is an example:
$\mathcal{M}_{3.3.A}:\left\{\begin{array}[]{ll}\dot{x}_{1}=-x_{1}\cdot x_{2}+p_{2}\cdot(10-x_{2}),\\ \dot{x}_{2}=-x_{1}\cdot x_{2}+(p_{2}+p_{3})\cdot(10-x_{2}),\\ \dot{x}_{3}=-p_{1}\cdot x_{3}+p_{3}\cdot(10-x_{2})\\ y_{1}=x_{1},y_{2}=x_{3}\end{array}\right.$ (9)
@@
This issue is not rare, and it is typically linked to a lack of accessibility from the problematic initial conditions. (Accessibility, or reachability, is the possibility of moving from an initial point in parameter space to a set with a non-empty interior.) Saccomani et al. (2003) studied it using a differential algebra approach.
The ORC can be computed from specific initial conditions: one can simply replace the symbolic state variables in the matrix, $x_{i},$ with specific numerical values. In this way, it is possible to test whether the model remains SLI from a given state. However, in doing so the stop criterion of the ORC is lost: it may no longer be sufficient to build $\mathcal{O}_{I}$ with $n_{x}+n_{\theta}-1$ Lie derivatives, as in (7). The following case study [^1] We thank Gleb Pogudin for noting this issue and providing the counter-example. illustrates this fact.
$\mathcal{M}_{3.3.B}:\left\{\begin{array}[]{ll}\dot{x}=1+\theta\cdot x^{2},\\ x(0)=0,\\ y=x.\end{array}\right.$ (10)

### v2

# Employing observability rank conditions for taking into account experimental information a priori

@@ Abstract
The concept of identifiability describes the possibility of inferring the parameters of a dynamic model by observing its output. It is common and useful to distinguish between structural and practical identifiability. The former property is fully determined by the model equations, while the latter is also influenced by the characteristics of the available experimental data. Structural identifiability can be determined by means of symbolic computations, which may be performed before collecting experimental data, and are hence sometimes called a priori analyses. Practical identifiability is typically assessed numerically, with methods that require simulations – and often also optimization – and are applied a posteriori. An approach to study structural local identifiability is to consider it as a particular case of observability, which is the possibility of inferring the internal state of a system from its output. Thus, both properties can be analysed jointly, by building a generalized observability matrix and computing its rank. The aim of this paper is to investigate to which extent such observability-based methods can also inform about practical aspects related with the experimental setup, which are usually not approached in this way. To this end, we explore a number of possible extensions of the rank tests, and discuss the purposes for which they can be informative as well as others for which they cannot.
@@ 1 Introduction
Identifiability is a broad concept related to mathematical modelling that is used in several contexts, from statistics to econometrics to dynamical systems. It refers to the possibility of inferring the parameters of a model by observing its output. In the area of system identification it is common to use the term structural identifiability. This concept was introduced by Bellman and Åström (1970), who used the word “structure” to refer to the equations of a dynamic model. This property determines whether it is possible “to get insight into the internal structure of a system from input-output measurements”. When analysing structural identifiability, one imposes no restrictions on the quality and quantity of data that can be obtained from the experiments. Thus, if a certain state variable, or function, is defined as a model output, the implicit assumption is that it can be sampled at an arbitrarily high rate and with absolute precision. Since this is an ideal scenario, another concept – practical identifiability – is needed to account for the additional uncertainties caused by the limitations of real-world data, which are noisy and scarce to a varying degree (Lam et al., 2022). Structural identifiability analyses provide a binary answer: a model is either structurally identifiable or unidentifiable. In contrast, practical identifiability analyses yield a numerical answer, typically in the form of a confidence interval in a parameter value. Overviews of practical and structural identifiability have been written by Miao et al. (2011) and more recently by Wieland et al. (2021).
@@
The observability of a system describes the possibility of inferring its internal state by observing its output. Here, internal state refers to the values at a given time of the set of time-varying variables given by the differential equations of the model. The concept was introduced by R. E. Kalman in the context of linear control systems. For nonlinear systems, Hermann and Krener (1977) presented an observability test called observability rank condition; Tunali and Tarn (1987) applied a similar condition to the study of identifiability. It should be noted that, if a parameter in a model is considered as a constant state variable, (structural local) identifiability can be studied in the same way as observability. Both concepts were reviewed jointly in (Villaverde, 2019).
@@ 2.1 Structural identifiability and observability
We say that a parameter $\theta_{i}\in\theta$ of $M$ (1) is structurally locally identifiable (SLI) if, for almost any vector $\theta^{*}\in\mathbb{R}^{n_{\theta}}$, there is a neighbourhood $\mathcal{N}(\theta^{*})$ such that, for any $\hat{\theta}\in\mathcal{N}(\theta^{*})$, $y(t,\theta^{*})=y(t,\hat{\theta})\iff\theta^{*}_{i}=\hat{\theta}_{i}$ (Walter and Pronzato, 1997; Chis et al., 2011; DiStefano, 2015). If this relationship does not hold in any $\mathcal{N}(\theta^{*})$, $\theta_{i}$ is structurally unidentifiable (SU). The phrase “almost any vector” means that the only possible exceptions belong to a set of measure zero.
The above definition entails that, if a parameter is SLI, its value can be determined from knowledge of the output $y(t)$ and input $u(t)$ vectors over time. Structural local identifiability is similar to another property, observability (Chatzis et al., 2015). Conceptually, a state variable $x_{i}(\tau)$ is observable if it can be determined from the output $y(t)$ and any known inputs $u(t)$ of the model in the interval $t_{0}\leq\tau\leq t\leq t_{f}$, for a finite $t_{f}$. Otherwise, it is unobservable. A model is called SLI if all its parameters are SLI, otherwise it is said to be SU. Likewise, a model is called observable if all its states are observable, and unobservable if at least one of them is unobservable (Wieland et al., 2021).
It should be noted that the properties defined above are structural but local, i.e. they tell us whether a parameter or a state variable can be uniquely estimated in a neighbourhood of (almost) any particular value, but they do not guarantee uniqueness in the whole of $\mathbb{R}^{n_{x}}$ or $\mathbb{R}^{n_{\theta}}.$ In contrast, a parameter is said to be structurally globally identifiable (SGI) if, for any $\hat{\theta}\in\mathcal{N}(\theta^{*})$ and almost any vector $\theta^{*}\in\mathbb{R}^{n_{\theta}}$, $y(t,\theta^{*})=y(t,\hat{\theta})\iff\theta^{*}_{i}=\hat{\theta}_{i}$, i.e. the neighbourhood $\mathcal{N}(\theta^{*})$ is the whole parameter space. A SGI parameter is SLI; a SLI parameter may be SGI (Walter and Pronzato, 1997).
@@ 2.2 A rank test for assessing observability
$M:\left\{\begin{aligned} \dot{x}&=&f\left(x,u\right),\\ y&=&g\left(x,u\right).\end{aligned}\right.$ (2)
Such a model is similar to the one given by (1), with the difference that it does not contain parameters (we will include them in the next subsection). There are several ways of analysing its observability. The core intuition that underlies some of them can be summarized as follows: first, define an object that contains the information that can be directly obtained from measuring the output, i.e., the output itself $y$ and its derivatives, $\dot{y},\ddot{y},...$. Then, establish a link between this mapping and the internal state of the system. Lastly, determine whether this mapping is unique, which would mean that the internal state can be derived from the output information.
A way of performing this test is to use the Taylor series expansion of the output function around the initial condition $x(t_{0})$, and then check whether this expansion is unique. This idea was proposed by Pohjanpalo (1978) for the purpose of structural identifiability – not observability – analysis. To perform this analysis, one must first obtain a set of algebraic equations on the parameters. If their solution is unique, then the parameters are SGI. A conceptually similar approach is the Generating Series introduced by Walter and Lecourtier (1982). Here, the output is expanded in series with respect to time and the inputs, whose coefficients are the output functions and their Lie derivatives. The Lie derivative $\mathcal{L}_{f}g(x,u)$ of the output function $g$ along the vector field $f$ is defined as:
$\mathcal{L}_{f}g(x,u)=\frac{\partial g(x,u)}{\partial x}f(x,u)+\sum_{j=0}^{j=\infty}\frac{\partial g(x,u)}{\partial u^{(j)}}u^{(j+1)}.$
The Lie derivative quantifies the rate of change of a vector field along the flow described by another vector field. In this case, it describes the change of the model output function $g$ along the dynamics of the system $f$, which is equal to the derivative of the output, $\dot{y}.$ Subsequent derivatives of the output, $\ddot{y},...,y^{(n_{x}-1)},$ are calculated as higher order Lie derivatives, which can be computed recursively from the lower order ones as follows (Karlsson et al., 2012):
$\mathcal{L}_{f}^{i}g(x,u)=\frac{\partial{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{L}_{f}^{i-1}}g(x,u)}{\partial x}f(x,u)+\sum_{j=0}^{j=\infty}\frac{\partial{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{L}_{f}^{i-1}}g(x,u)}{\partial u^{(j)}}u^{(j+1)}$ (3)
In the generating series approach, one builds an exhaustive summary that consists of the coefficients of the output and their successive Lie derivatives up to some order, evaluated at the initial condition $x(t_{0})$. If the exhaustive summary is unique, then the parameters are SGI.
Here we adopt an approach that has similarities to the Taylor and Generating Series methods outlined above. It is sometimes called the geometric approach, and it follows the line of work initiated in the 1970s by a number of researchers, including Griffith and Kumar (1971); Kou et al. (1973); Hermann and Krener (1977). It begins by defining an “observability mapping” that consists of the output and its time derivatives, and it is obtained by calculating $n_{x}-1$ Lie derivatives of $g(x,u)$. Then, the relation between the observability mapping and the internal state of the system is obtained by computing an “observability matrix”, $\mathcal{O}(x,u)$. This matrix consists of the Jacobian of the observability mapping with respect to the state variables, i.e.:
@@
A note about the calculation of the Lie derivatives is in order. Note that the expressions in (3) contain infinite summations. However, since $\dot{u}$ does not appear in the model equations (1), the summations can be truncated so that (Villaverde et al., 2019):
$\mathcal{L}_{f}g(x,u)$ $=\frac{\partial g(x,u)}{\partial x}f(x,u),$ (5)
@@ 2.4 Practical Identifiability
3 Modifying the rank test to take into account experimental information
In this section we investigate in which ways we can use structural analyses – and more specifically the observability-based approach described in Sections 2.12.3 – to obtain information about identifiability for specific experimental configurations. Thus, our investigation will be centred around one object, the observability-identifiability matrix $\mathcal{O}_{I}$ defined in (8), and we will discuss a number of directions in which to extend it or modify it so as to take into account ‘practical’ aspects of the estimation process.
@@ 3.1 Multiple experiments
Ovchinnikov et al. (2022) presented an algorithm that computes (i) the smallest number of experiments to maximise local identifiability, and (ii) the number of experiments to maximise global identifiability so that this number exceeds the minimal such number by at most one. This algorithm combines a differential algebra approach with observability rank checks.
@@ 3.2 Input characteristics
A question related to Sections 3.1 and 3.2 concerns the implementation of inputs in real applications, and how to modify the analyses to approximate them. For example, instead of applying an ideal ramp, due to experimental constraints we may be forced to apply instead a piecewise constant input. How should we treat this scenario in our analyses? One the one hand, an input consisting of a series of constant values resembles a set of multiple experiments with constant inputs. On the other hand, as the number of steps increases the input becomes closer to an actual ramp. Wang and Sontag (1989) showed that the observation spaces defined by means of piecewise constant inputs are equivalent to those defined by differentiable inputs. This result supports the idea of adopting the latter of the two aforementioned views–albeit with the caveats that usually apply to the extrapolation of results from ideal experiments to real ones. The following example illustrates this case.
@@
Let us now perform a numerical experiment: we simulate the model to generate artificial data, applying as input a piecewise constant signal. Since our focus is on the role played by the input, we add a small amount of noise to the output, in order to yield a realistic experiment while avoiding confounding effects. We generate 100 realizations of Gaussian noise with zero mean and $\sigma=0.1$. Next, we estimate the parameters by fitting the model to each realization of the data, using a hybrid metaheuristic optimization algorithm (Egea et al., 2014). Lastly, we use a bootstrap procedure to quantify their uncertainty (Balsa-Canto et al., 2010). The left panel of Figure 1 shows one realization of the pseudo-experimental data, along with the model states, and the input. The right hand panel of Figure 1 shows the optimization results for one of the parameters, $k_{12}$. It can be seen that, while there is some uncertainty associated with the estimate, both the mean and median estimations are very close to the nominal value ($k_{12}=7$), and the confidence intervals are reasonably narrow. It is clear from these results that $k_{12}$ is identifiable. Hence, performing the observability rank test by approximating a piecewise constant input as a ramp yields correct SIA results (at least in this case study), while the approximation as a multi-experiment with constant inputs yields a wrong result. Intuitively, this can be understood as a failure of the multi-experiment setup to capture the continuity between experiments; a set of independent experiments does not provide all the information.
Figure 1: Model $\mathcal{M}_{3.2}$. The panel on the left shows simulations of the model input, $u(t),$ and the output $y(t)=x_{1}(t);$ the blue circles represent the artificial noiseless data used for parameter estimation. The panel on the right shows the resulting bootstrap of parameter $k_{12}$.
* Unknown inputs
A related issue of practical importance concerns unknown inputs. So far we have assumed that the inputs $u(t)$ are perfectly known, but in reality we may face challenges such as non-negligible dynamic disturbances, unknown parameters that are time-varying, or external inputs that are known only to a certain extent. These problems can be seen as different ways of approaching the existence of unknown inputs. They do not only make estimation more difficult, but also complicate the observability analyses.
In recent years, a number of results have been published that extend the ORC to enable the inclusion of unknown inputs (Martinelli, 2022; Maes et al., 2019). These works consider the unknown inputs and their derivatives, up to a given order, as additional state variables. The idea of limiting the number of nonzero input derivatives is similar to the assumption discussed for known inputs, although the computational cost is higher in the unknown input case. An efficient algorithm to check observability of models with unknown inputs was presented by Shi and Chatzis (2022).
@@ 3.3 Initial conditions
The definitions of structural local identifiability and observability are valid ‘almost everywhere’, that is, with the possible exception of a set of measure zero. Thus, a model that is SLI may become unidentifiable for specific initial conditions. The simplest example of this phenomenon is the model given by $\dot{x}=p\cdot x,$ $y=x.$ It is obvious that $x$ is observable and $p$ is SLI, except if $x(0)=0,$ in which case the system remains at the initial state and it is not possible to infer $p.$ Let us check the ORC for this model. Augmenting the state as explained in Section 2.3, we obtain
$\mathcal{M}_{3.3.A}:\left\{\begin{array}[]{ll}\dot{x}_{1}=x_{1}\cdot x_{2},\\ \dot{x}_{2}=0,\\ y_{1}=x_{1}.\end{array}\right.$ (10)
Since the model has two state variables, one Lie derivative suffices to obtain the observability-identifiability matrix as follows:
$\mathcal{O}_{I}=\begin{pmatrix}1&0\\ x_{2}&x_{1}\end{pmatrix}$ (11)
The rank of this matrix is two, except for $x_{1}=0.$ Thus, the ORC result is in agreement with the loss of identifiability from the specific initial condition. This issue is not rare, and it is typically linked to a lack of accessibility from the problematic initial conditions. (Accessibility, or reachability, is the possibility of moving from an initial point in parameter space to a set with a non-empty interior.) Saccomani et al. (2003) studied it using a differential algebra approach. The following model (August and Papachristodoulou, 2009) provides a more realistic example:
$\mathcal{M}_{3.3.B}:\left\{\begin{array}[]{ll}\dot{x}_{1}=-x_{1}\cdot x_{2}+p_{2}\cdot(10-x_{2}),\\ \dot{x}_{2}=-x_{1}\cdot x_{2}+(p_{2}+p_{3})\cdot(10-x_{2}),\\ \dot{x}_{3}=-p_{1}\cdot x_{3}+p_{3}\cdot(10-x_{2}),\\ y_{1}=x_{1},y_{2}=x_{3}.\end{array}\right.$ (12)
@@
An important warning should be made here. In the example (10), we computed the ORC from a given initial condition simply by replacing the symbolic state variables in the matrix, $x_{i},$ with specific numerical values. In this way, it is possible to test whether the model remains SLI from a given state. However, in doing so the stop criterion of the ORC is lost: it may no longer be sufficient to build $\mathcal{O}_{I}$ with $n_{x}+n_{\theta}-1$ Lie derivatives, as in (8). The following case study [^1] We thank Gleb Pogudin for noting this issue and providing the counter-example. illustrates this fact.
$\mathcal{M}_{3.3.C}:\left\{\begin{array}[]{ll}\dot{x}=1+\theta\cdot x^{2},\\ x(0)=0,\\ y=x.\end{array}\right.$ (13)